OpenStudy (anonymous):

How do you solve 3|5x-1|+9(less than or equal to) 23

OpenStudy (jdoe0001):

by \(\bf \mathbb{NIKE}\) :P

OpenStudy (jdoe0001):

you'd do it the same way you'd any equation more or less so \(\bf 3|5x-1|+9 \le 23\) you proceed about the same as you'd with \(\bf 3|5x-1|+9 = 23\)

OpenStudy (jdoe0001):

just keep in mind that |5x-1| is really meant to be +1(5x-1) AND -1(5x-1) so \(\bf 3|5x-1|+9 \le 23\\ \color{blue}{3\left[1(5x-1)\right]\le 23\\ 3\left[-1(5x-1)\right]\le 23}\)

OpenStudy (anonymous):

I feel stupid...thank you

OpenStudy (jdoe0001):

yw

OpenStudy (jdoe0001):

the one exception you want to keep in mind is that, in inequalities whenever you multiply or divide or exponentialize by a negative value you \(\bf \color{red}{\text{ flip the sign}}\) only when you use a negative factor though

OpenStudy (jdoe0001):

so say -2x < 10 # multiplying both sides by -2 x > -5

OpenStudy (jdoe0001):

well, multiplying by - 1/2 rather